A temperature difference of 5 K is equal to
1
Expert's answer
2020-04-20T10:12:58-0400

A temperature difference of 5 K is equal to a difference of 5 on the Celsius scale. Also it will be a difference of 9 on the Fahrenheit scale.


"\Delta T_F=\frac{9}{5}\Delta T_C=\frac{9}{5}\Delta T_K=\frac{9}{5}5=9\degree F"